# Uncertainty Quantification of Pre-Trained and Fine-Tuned Surrogate Models using Conformal Prediction

### v1

@@ 2.3 Exchangeablity Assumptions
Within the experiments considered within this paper, surrogate modelling of spatio-temporal data is treated as an initial value problem (IVP). To map the evolution of the system under study, the surrogate model takes in the initial state(s) of the system and evolves the system autoregressively in time. By treating the modelling task as an initial value problem, each input-output pair ($X_{i},Y_{i}$) from the calibration and the prediction dataset is assumed to be exchangeable. The initial condition associated with each IVP is sampled i.i.d. from a distribution characterising the possible (large) range of interested initial conditions for the prediction task. The assumption relies on the further caveat that as outlined in Section 2.1.3, the structure of each spatio-temporal output is maintained. The CP framework requires exchangeability not just across the calibration data but across the prediction data as well (Angelopoulos and Bates, 2023). To maintain exchangeability across the calibration and prediction datasets, we have to ensure that the full spatio-temporal context window across them remains the same and assume that they arise from the same distribution. The exchangeability assumption is more complex in the case of experimental data, and we discuss this further within the experiments in Section 3.7 and Section 3.8.
@@
Within Section 3 we consider two types of experiments: surrogate modelling over simulations and experimental data. Within the simulation context (Section 3.1 to Section 3.6), the modelling task takes in the initial state of the system at $t=0$ and then evolves the system (autoregressively or in one-shot) to time $t=T$. The model always starts at the initial condition(s) of the simulation and maps till a fixed time state $T$. Whereas in those modelling experimental data (Section 3.7 and Section 3.8) the modelling task takes the system at $t=T$ to $t=T+\Delta T$. The starting time instance (initial condition) $T$ changes in each exchangeable data pair but the model evolves it for a fixed time duration of $\Delta T$. Within the simulation setup, it is easier to gather data as required, exploring a data-abundant scenario. In the case of real-world experiments, data is scarce and we had to rely on breaking down available data to obtain the calibration data. We characterise this difference in how the exchangeable pairs are conceived for both the simulation and experimental data in Figure 4.
@@ 3 Experiments
To evolve spatio-temporal data describing physical systems, we train several surrogate models using simulation data built from numerical solvers designed for the PDE of interest. Across our experiments, we choose different kinds of surrogate models. Multi-Layer Perceptrons (MLPs) (Haykin, 1994), U-Nets (Ronneberger et al., 2015), Fourier Neural Operators (FNOs) (Li et al., 2021), Vision Transformers (Yin et al., 2022b), and Graph Neural Networks Scarselli et al. (2009) are utilised. These networks are commonly used as surrogate models within physics and engineering disciplines (Mánek et al., 2023; Gupta and Brandstetter, 2023; Wen et al., 2023; Geneva and Zabaras, 2022; Sanchez-Gonzalez et al., 2020).
MLPs have been used as surrogates for modelling wind turbine blades (Lalonde et al., 2021), classifiers for high energy physics (Baldi et al., 2016), and for designing fusion reactors (Mánek et al., 2023). U-Nets and FNOs are structured as neural-PDE based approximators that can learn across the operator space rather than the function space of the mapping. They find utility in designing surrogates for fluid dynamics (Gupta and Brandstetter, 2023), studying carbon capture and mitigation (Wen et al., 2023), weather modelling (Kurth et al., 2023) and modelling plasma evolution within a nuclear fusion device (Gopakumar et al., 2023). Transformers have found application in PDE solving (Li et al., 2023b) and weather modelling (Nguyen et al., 2023). GNNs have been deployed to model Lagrangian dynamics of a PDE and in solving particle-in-cell scenarios (Brandstetter et al., 2022); they also have been applied to weather forecasting (Lam et al., 2023).
The MLP represents a fully connected neural network. The U-Net is a fully convolutional autoencoder (Hinton and Salakhutdinov, 2006) with skip connections (He et al., 2016; Takamoto et al., 2022). The FNO is an operator learning method that learns the kernel integration across the Fourier space (Li et al., 2021). Vision Transformers is an attention-based model that learns spatio-temporal rollouts through sequence patching of the spatial outputs and auto-regressive rollouts. Graph Neural Networks deploy message-passing across nodes representing the spatio-temporal information to model the evolution of a physical system. Depending on the specific experiment we adjust the model setup to either do a one-step forward-in-time mapping or perform an auto-regressive time roll-out over multiple time steps. All models were trained on an NVIDIA A100 GPU, while the calibration and estimation of prediction sets were done on a standard laptop.
@@
Table 1 forms the centrepiece of our work. The table shows the coverage obtained when predictions are calibrated for $\alpha=0.1$ (90% coverage) across our experiments utilising the three different calibration methods (non-conformity scores) discussed in Section 2.2. For CQR and STD, we additionally show the coverage before performing CP, as these methods have initial uncertainty estimates, which are calibrated to ensure guaranteed coverage. This is labelled as ‘uncalibrated’ in Table 1. For each experiment, an adequate pre-trained surrogate model is developed or chosen. The CP framework is performed over its output dimensions, observing near-to-ideal coverage. The “tightness of fit” depends on the choice of non-conformity score. The tightness is taken to be the width of the calibrated error bars, averaged across each cell of the tensor. In most experiments we find the AER to offer the best fit, however, STD is a close second.
@@
The Poisson equation is a generalisation of the Laplace Equation. It relates the Laplacian of a field to the distribution of the sources/sinks in the field (Hackbusch, 2017). The elliptic PDE models various physical phenomena, such as electrostatics, gravitation and fluid dynamics. The Poisson equation is a steady-state problem (spatial-only), where the interest lies in mapping the initial distribution of the field to its final state. The state is given along the x-axis, discretising the domain $[0,1]$ into 32 uniform grid points. An initial dataset consisting of 5000 simulation data points is generated and used to train an MLP. Another 1000 data points are generated to calibrate the trained model using the CP framework along with another 1000 for validating the procedure. The training data, calibration, and validation datasets are sampled from the same distribution within this case. Further details about the physics, data generation strategies and model training can be found in Appendix A.
@@
Multiple MLPs are trained to demonstrate the efficacy of different non-conformity scores while using the CP framework. To perform the CQR, three models are trained, modelling the $5^{th}$, $50^{th}$ and $95^{th}$ quantile, while for AER a single model is trained to output the evolved field using an L1-norm. For CP using STD, an MLP with dropout is trained to obtain a probabilistic model that learns the distribution of the evolved state. The performance comparison of each non-conformity score is laid out in Table 1.
@@ 3.3.1 U-Net
Figure 7: Calibrated prediction sets obtained by performing conformal prediction with the three non-conformity scores mentioned in Section 2.2 for the 2D wave equation. The color bar depicts the coverage (in %) provided by each prediction set, characterised by the $1-\alpha$ value. Each figure shows a slice of the prediction set and exact inference value taken along the $y$-axis for the final time instance of the prediction. The exact numerical solution (black with white markers) can be seen to fit up to the coverage offered by the prediction set.
@@ 3.8 Camera Diagnostic on a Tokamak
In our previous work (Gopakumar et al., 2024), we demonstrated the utility of the FNO to forecast the evolution of the plasma within the Tokamak. Within that work, the FNO was trained to take in an input sequence of 10 camera frames and predict a sequence of the next 10 camera frames of the plasma within the Tokamak. The model is trained on sequences of frames characterising the entire duration of a plasma shot within the MAST Tokamak. In this work, we apply the CP framework across the FNO-camera using the AER non-conformity score to provide valid error bars to the plasma evolution forecasts. For results and further information about the camera data, FNO and the training setup used, we advise the reader to refer to the original work by (Gopakumar et al., 2024).
@@
We have the following arguments to support the validity of using CP for this time dependant problem, 1) only the calibration dataset and the prediction $X_{n+1}$ are required to be exchangeable, 2) we predict errors over an entire time interval simultaneously, not sequentially, and 3) we assume there is no large shift in the data between the calibration shot and prediction shot. The exchangeability requirement across the calibration and prediction shot requires them to follow similar plasma discharge profiles and device conditions. Predicting across a characteristically different shot violates this assumption and suffers from lack of coverage as shown in Appendix I.
@@ 4.2 Limitations
Input Independence. Uncertainty ($\sim\hat{q}$) is added to the model outputs as prediction sets are independent of the model inputs and a function of the calibration dataset. In more Bayesian methods, the prediction uncertainty, estimated in the posterior is seen as a function of the prior and the likelihood, hence varying for each region of the prediction regime. Within the deterministic models explored within this work (using non-conformity score AER), the value of the error ($\sim\hat{q}$) is determined before the model prediction and remains fixed across the inference regime. This is exceptionally clear in the case of the fine-tuned foundation physics model (see Figure 12, where the error bars indicate the global error approximation and not on a case-by-case basis. For probabilistic models (using STD non-conformity score) this is less severe as the error bar is not just a function of $\hat{q}$ but also of the standard deviation output from the model. This allows for the prediction sets to be more flexible and adjust as a function of the model inputs (see Figure 2). Though there has been work on making the CP framework further input dependent (Johansson et al., 2021), they still require augmenting deterministic models through ensembles or architectural changes.
@@ 5 Conclusion
In this paper, it is shown that by maintaining the exchangeability of the data across the spatio-temporal domain (i.e. by preserving the tensorial shape), conformal prediction provides valid error bars (i.e., satisfied Equation 1 to expensive surrogate models for scientific applications, even if they make out-of-distribution predictions. Our empirical studies demonstrate that conformal prediction can be applied to any pre-trained / fine-tuned model for obtaining statistically guaranteed coverage, irrespective of the model configuration, training conditions and dimensionality of the outputs. One of the crucial highlights of the study is that CP can help provide guaranteed prediction sets across a new set of PDE solutions (out-of-distribution data), different to that which the model was conditioned on. Our work benchmarks against three kinds of calibration methods (non-conformity scores), across deterministic and probabilistic models, ensuring validity for the coverage. Although we focus on providing guaranteed coverage for modelling spatio-temporal data, we believe that our findings extend to any models predicting outputs with fixes fixed tensorial output shapes and with exchangeable calibration and prediction regimes.
@@ Appendix A Poisson Equation
A total of 7,000 data points are generated, where 5,000 are used to train an MLP with 3 layers and 64 neurons in each layer, 1,000 are used to perform the calibration required to estimate the non-conformity scores and another 1,000 for validation. Being a steady-state problem, the MLP learns how a scalar field evolves under the influence of the Laplacian, mapping from the initial to the final state of evolution. The network learns to map the initial condition to the final steady-state solution.
@@ D.2 Model and Training
We train an FNO to map the spatio-temporal evolution of the vorticity taking in the first 10 time instances (T_in) to the next 10 time instances (step). For the case of the Navier-Stokes equations, we deploy a feed-forward mapping from the initial 10 time steps to the next 10 time steps. The architecture for the FNO can be found in Table 8. We deploy a Min-Max normalisation strategy, allowing the field values to lie between -1 and 1. Each model is trained for up to 500 epochs using the Adam optimiser (Kingma and Ba, 2015) with a step decaying learning rate. The learning rate is initially set to 0.005 and scheduled to decrease by half after every 100 epochs. The model was trained using a relative LP loss. Considering the efficiency and simplicity, conformal prediction for the Navier-Stokes case was conducted using AER and STD as a non-conformity metric as given in the Section 2.2.
@@ E.2 Model and Training
We train a multi-variable FNO to map the spatio-temporal evolution of the field variable taking in the first 10 time instances (T_in) to the next 5 time instances (step). For the case of the MHD equations, we deploy an auto-regressive structure that performs a time rollout allowing us to map the initial time steps in a recursive manner up until the desired time instance (T_out). Each model autoregressively models the evolution of the field variable up until the $50^{th}$ time instance. The architecture for the multi-variable FNO can be found in Table 10. We deploy a two-fold normalisation strategy considering the nature of the dataset. The physical field information represented within the MHD cases are in different scales, with densities ranging from 0 to $201e$ 20 and temperatures ranging up to $61e$ 6. Since we are considering the gradual diffusion of an inhomogeneous density blob(s), the data distribution within the spatial domain is severely imbalanced. Taking these aspects of the training data into consideration, a physics normalisation is performed initially, where the field values are scaled down by dividing against the prominent field value. This is followed up by a linear range scaling, allowing the field values to lie between -1 and 1. Each model is trained for up to 500 epochs using the Adam optimiser (Kingma and Ba, 2015) with a step decaying learning rate. The learning rate is initially set to 0.005 and scheduled to decrease by half after every 100 epochs. The model was trained using a relative LP loss. Considering the efficiency and simplicity, conformal prediction for the MHD case was only conducted using the AER as a non-conformity metric as given in the Section 2.2.
@@ Appendix G Impact of Calibration Size
We conduct an empirical study exploring the impact the size of the calibration dataset has on providing guaranteed coverage within our experiments. We iterate over $n_{cal}=250,500,750,1000$ for the Poisson (Figure 28), Convection-Diffusion (Figure 29) and the Wave Equation (Figure 30, 31). Though the coverage obtained is from a Beta distribution governed by $n_{cal}$ and changes with each sampling from that, our experiments are restricted to a single sample of $n_{cal}$ data points from that distribution. The study of the impact of the calibration dataset is done across all the various non-conformity scores as well.
Figure 28: Impact of the size of the calibration dataset on the coverage obtained for Conformal Prediction across various non-conformity scores for the Poisson Equation. Irrespective of the chosen size of the calibration dataset, we obtain guaranteed coverage.
Figure 29: Impact of the size of the calibration dataset on the coverage obtained for Conformal Prediction across various non-conformity scores for the Convection-Diffusion Equation. Irrespective of the chosen size of the calibration dataset, we obtain guaranteed coverage, however with larger $n_{cal}$ we obtain marginally better coverage. Irrespective of the chosen size of the calibration dataset, we obtain guaranteed coverage.
Figure 30: Impact of the size of the calibration dataset on the coverage obtained for Conformal Prediction across various non-conformity scores for the Wave Equation modelled using U-Net. Irrespective of the chosen size of the calibration dataset, we obtain guaranteed coverage.
Figure 31: Impact of the size of the calibration dataset on the coverage obtained for Conformal Prediction across various non-conformity scores for the Wave Equation modelled using FNO. Irrespective of the chosen size of the calibration dataset, we obtain guaranteed coverage.
From Figure 28 to Figure 31, we explore the impact of the size of the calibration dataset on the coverage obtained for the various non-conformity scores. We notice that across our experiments, we obtain guaranteed coverage, irrespective of the chosen non-conformity score or the size of the calibration dataset ($n_{cal}$). Though the size of the calibration dataset has no impact on the guarantee as can be witnessed in the above figures, the tightness of the error bars is governed by the size of the calibration dataset.

### v2

@@ 2.3 Exchangeablity Assumptions
Throughout the experiments considered in this paper, surrogate modelling of spatio-temporal data is treated as an initial value problem (IVP). To map the evolution of the system under study, the surrogate model takes in the initial state(s) of the system and evolves the system autoregressively in time. By treating the modelling task as an initial value problem, each input-output pair ($X_{i},Y_{i}$) from the calibration and the prediction dataset is assumed to be exchangeable. The initial condition associated with each IVP is sampled i.i.d. from a distribution characterising the possible (large) range of interested initial conditions for the prediction task. The assumption relies on the further caveat that is outlined in Section 2.1.2, the structure of each spatio-temporal output is maintained. The CP framework requires exchangeability not just across the calibration data but across the prediction data as well (Angelopoulos and Bates, 2023). To maintain exchangeability across the calibration and prediction datasets, we have to ensure that the full spatio-temporal context window across them remains the same and assume that they arise from the same distribution. The exchangeability assumption is more complex in the case of experimental data, and we discuss this further in Section 3.7 and Section 3.8.
@@
In Section 3 we consider two types of experiments: surrogate modelling over simulations and experimental data. Within the simulation context (Section 3.1 to Section 3.6), the modelling task takes in the initial state of the system at $t=0$ and then evolves the system (autoregressively or in one-shot) to time $t=T$. The model always starts at the initial condition(s) of the simulation and maps till a fixed time state $T$. Whereas in those modelling experimental data (Section 3.7 and Section 3.8) the modelling task takes the system at $t=T$ to $t=T+\Delta T$. The starting time instance (initial condition) $T$ changes in each exchangeable data pair but the model evolves it for a fixed time duration of $\Delta T$. Within the simulation setup, it is easier to gather data as required, exploring a data-abundant scenario. In the case of real-world experiments, data is scarce and we had to rely on breaking down available data to obtain the calibration data. We characterise this difference in how the exchangeable pairs are conceived for both the simulation and experimental data in Figure 4.
@@ 3 Experiments
To evolve spatio-temporal data describing physical systems, we train several surrogate models using simulation data built from numerical solvers designed for the PDE of interest. Across our experiments, we choose different kinds of surrogate models. Multi-Layer Perceptrons (MLPs) (Haykin, 1994), U-Nets (Ronneberger et al., 2015), Fourier Neural Operators (FNOs) (Li et al., 2021), vision transformers (Yin et al., 2022b), and graph neural networks Scarselli et al. (2009) are utilised. These networks are commonly used as surrogate models for challenging systems in physics and engineering (Mánek et al., 2023; Gupta and Brandstetter, 2023; Wen et al., 2023; Geneva and Zabaras, 2022; Sanchez-Gonzalez et al., 2020).
MLPs have been used as surrogates for modelling wind turbine blades (Lalonde et al., 2021), classifiers for high-energy physics (Baldi et al., 2016), and for designing fusion reactors (Mánek et al., 2023). U-Nets and FNOs are structured as neural-PDE based approximators that can learn across the operator space rather than the function space of the mapping. They find utility in designing surrogates for fluid dynamics (Gupta and Brandstetter, 2023), studying carbon capture and mitigation (Wen et al., 2023), weather modelling (Kurth et al., 2023) and modelling plasma evolution within a nuclear fusion device (Gopakumar et al., 2023). Transformers have found application in PDE solving (Li et al., 2023b) and weather modelling (Nguyen et al., 2023). GNNs have been deployed to model Lagrangian dynamics of a PDE and in solving particle-in-cell scenarios (Brandstetter et al., 2022); they also have been applied to weather forecasting (Lam et al., 2023).
The MLP represents a fully connected neural network. The U-Net is a fully convolutional autoencoder (Hinton and Salakhutdinov, 2006) with skip connections (He et al., 2016; Takamoto et al., 2022). The FNO is an operator learning method that learns the kernel integration across the Fourier space (Li et al., 2021). A vision transformer is an attention-based model that learns spatio-temporal rollouts through sequence patching of the spatial outputs and auto-regressive rollouts. Graph neural networks deploy message-passing across nodes representing the spatio-temporal information to model the evolution of a physical system. Depending on the specific experiment we adjust the model setup to either do a one-step forward-in-time mapping or perform an auto-regressive time roll-out over multiple time steps. All models were trained on an NVIDIA A100 GPU, while the calibration and estimation of prediction sets were done on a standard laptop.
@@
Table 1 forms the centrepiece of our work. The table shows the coverage obtained when predictions are calibrated for $\alpha=0.1$ (90% coverage) across our experiments utilising the three different calibration methods (nonconformity scores) discussed in Section 2.2. For CQR and STD, we additionally show the coverage before performing CP, as these methods have initial uncertainty estimates, which are calibrated to ensure guaranteed coverage. This is labelled as ‘uncalibrated’ in Table 1. For each experiment, an adequate pre-trained surrogate model is developed or chosen. The CP framework is performed over its output dimensions, observing near-to-ideal coverage. The “tightness of fit” depends on the choice of nonconformity score. The tightness is taken to be the width of the calibrated error bars, averaged across each cell of the tensor. In most experiments we find the AER to offer the best fit, however, STD is a close second.
@@
The Poisson equation is a generalisation of the Laplace equation. It relates the Laplacian of a field to the distribution of the sources/sinks in the field (Hackbusch, 2017). The elliptic PDE models various physical phenomena, such as electrostatics, gravitation, and fluid dynamics. The Poisson equation is a steady-state problem (spatial-only), where the interest lies in mapping the initial distribution of the field to its final state. The state is given along the x-axis, discretising the domain $[0,1]$ into 32 uniform grid points. An initial dataset consisting of 5000 simulation data points is generated and used to train an MLP. Another 1000 data points are generated to calibrate the trained model using the CP framework along with another 1000 for validating the procedure. The training data, calibration, and validation datasets are sampled from the same distribution within this case. Further details about the physics, data generation strategies and model training can be found in Appendix A.
@@
Multiple MLPs are trained to demonstrate the efficacy of different nonconformity scores while using the CP framework. To perform the CQR, three models are trained, modelling the $5^{th}$, $50^{th}$ and $95^{th}$ quantile, while for AER a single model is trained to output the evolved field using an L1-norm. For CP using STD, an MLP with dropout is trained to obtain a probabilistic model that learns the distribution of the evolved state. The performance comparison of each nonconformity score is laid out in Table 1.
@@ 3.3.1 U-Net
Figure 7: Calibrated prediction sets obtained by performing conformal prediction with the three nonconformity scores mentioned in Section 2.2 for the 2D wave equation. The color bar depicts the coverage (in %) provided by each prediction set, characterised by the $1-\alpha$ value. Each figure shows a slice of the prediction set and exact inference value taken along the $y$-axis for the final time instance of the prediction. The exact numerical solution (black with white markers) can be seen to fit up to the coverage offered by the prediction set.
@@ 3.8 Camera Diagnostic on a Tokamak
In our previous work (Gopakumar et al., 2024), we demonstrated the utility of the FNO to forecast the evolution of the plasma within the Tokamak. Within that work, the FNO was trained to take in an input sequence of 10 camera frames and predict a sequence of the next 10 camera frames of the plasma within the Tokamak. The model is trained on sequences of frames characterising the entire duration of a plasma shot within the MAST Tokamak. In this work, we apply the CP framework across the FNO-camera using the AER nonconformity score to provide valid error bars to the plasma evolution forecasts. For results and further information about the camera data, FNO and the training setup used, we advise the reader to refer to the original work by (Gopakumar et al., 2024).
@@
We have the following arguments to support the validity of using CP for this time dependant problem, 1) only the calibration dataset and the prediction $X_{n+1}$ are required to be exchangeable, 2) we predict errors over an entire time interval simultaneously, not sequentially, and 3) we assume there is no large shift in the data between the calibration shot and prediction shot. The exchangeability requirement across the calibration and prediction shot requires them to follow similar plasma discharge profiles and device conditions. Predicting across a characteristically different shot violates this assumption and suffers from lack of coverage as shown in Figure 21.
@@ 4.2 Limitations
Input Independence. Uncertainty ($\sim\hat{q}$) is added to the model outputs as prediction sets are independent of the model inputs and a function of the calibration dataset. In more Bayesian methods, the prediction uncertainty, estimated in the posterior is seen as a function of the prior and the likelihood, hence varying for each region of the prediction regime. Within the deterministic models explored within this work (using nonconformity score AER), the value of the error ($\sim\hat{q}$) is determined before the model prediction and remains fixed across the inference regime. This is exceptionally clear in the case of the fine-tuned foundation physics model (see Figure 12, where the error bars indicate the global error approximation and not on a case-by-case basis. For probabilistic models (using STD nonconformity score) this is less severe as the error bar is not just a function of $\hat{q}$ but also of the standard deviation output from the model. This allows for the prediction sets to be more flexible and adjust as a function of the model inputs (see Figure 2). Though there has been work on making the CP framework further input dependent (Johansson et al., 2021), they still require augmenting deterministic models through ensembles or architectural changes.
@@ 5 Conclusion
In this paper, it is shown that by maintaining the exchangeability of the data across the spatio-temporal domain (i.e. by preserving the tensorial shape), conformal prediction provides valid error bars (i.e., satisfied Equation 1 to expensive surrogate models for scientific applications, even if they make out-of-distribution predictions. Our empirical studies demonstrate that conformal prediction can be applied to any pre-trained / fine-tuned model for obtaining statistically guaranteed coverage, irrespective of the model configuration, training conditions and dimensionality of the outputs. One of the crucial highlights of the study is that CP can help provide guaranteed prediction sets across a new set of PDE solutions (out-of-distribution data), different to that which the model was conditioned on. Our work benchmarks against three kinds of calibration methods (nonconformity scores), across deterministic and probabilistic models, ensuring validity for the coverage. Although we focus on providing guaranteed coverage for modelling spatio-temporal data, we believe that our findings extend to any models predicting outputs with fixes fixed tensorial output shapes and with exchangeable calibration and prediction regimes.
@@ Appendix A Poisson Equation
A total of 7,000 data points are generated, where 5,000 are used to train an MLP with 3 layers and 64 neurons in each layer, 1,000 are used to perform the calibration required to estimate the nonconformity scores and another 1,000 for validation. Being a steady-state problem, the MLP learns how a scalar field evolves under the influence of the Laplacian, mapping from the initial to the final state of evolution. The network learns to map the initial condition to the final steady-state solution.
@@ D.2 Model and Training
We train an FNO to map the spatio-temporal evolution of the vorticity taking in the first 10 time instances (T_in) to the next 10 time instances (step). For the case of the Navier-Stokes equations, we deploy a feed-forward mapping from the initial 10 time steps to the next 10 time steps. The architecture for the FNO can be found in Table 8. We deploy a Min-Max normalisation strategy, allowing the field values to lie between -1 and 1. Each model is trained for up to 500 epochs using the Adam optimiser (Kingma and Ba, 2015) with a step decaying learning rate. The learning rate is initially set to 0.005 and scheduled to decrease by half after every 100 epochs. The model was trained using a relative LP loss. Considering the efficiency and simplicity, conformal prediction for the Navier-Stokes case was conducted using AER and STD as a nonconformity metric as given in the Section 2.2.
@@ E.2 Model and Training
We train a multi-variable FNO to map the spatio-temporal evolution of the field variable taking in the first 10 time instances (T_in) to the next 5 time instances (step). For the case of the MHD equations, we deploy an auto-regressive structure that performs a time rollout allowing us to map the initial time steps in a recursive manner up until the desired time instance (T_out). Each model autoregressively models the evolution of the field variable up until the $50^{th}$ time instance. The architecture for the multi-variable FNO can be found in Table 10. We deploy a two-fold normalisation strategy considering the nature of the dataset. The physical field information represented within the MHD cases are in different scales, with densities ranging from 0 to $201e$ 20 and temperatures ranging up to $61e$ 6. Since we are considering the gradual diffusion of an inhomogeneous density blob(s), the data distribution within the spatial domain is severely imbalanced. Taking these aspects of the training data into consideration, a physics normalisation is performed initially, where the field values are scaled down by dividing against the prominent field value. This is followed up by a linear range scaling, allowing the field values to lie between -1 and 1. Each model is trained for up to 500 epochs using the Adam optimiser (Kingma and Ba, 2015) with a step decaying learning rate. The learning rate is initially set to 0.005 and scheduled to decrease by half after every 100 epochs. The model was trained using a relative LP loss. Considering the efficiency and simplicity, conformal prediction for the MHD case was only conducted using the AER as a nonconformity metric as given in the Section 2.2.
@@ Appendix G Impact of Calibration Size
We conduct an empirical study exploring the impact the size of the calibration dataset has on providing guaranteed coverage within our experiments. We iterate over $n_{cal}=250,500,750,1000$ for the Poisson (Figure 29), Convection-Diffusion (Figure 30) and the Wave Equation (Figure 31, 32). Though the coverage obtained is from a Beta distribution governed by $n_{cal}$ and changes with each sampling from that, our experiments are restricted to a single sample of $n_{cal}$ data points from that distribution. The study of the impact of the calibration dataset is done across all the various nonconformity scores as well.
Figure 29: Impact of the size of the calibration dataset on the coverage obtained for Conformal Prediction across various nonconformity scores for the Poisson Equation. Irrespective of the chosen size of the calibration dataset, we obtain guaranteed coverage.
Figure 30: Impact of the size of the calibration dataset on the coverage obtained for Conformal Prediction across various nonconformity scores for the Convection-Diffusion Equation. Irrespective of the chosen size of the calibration dataset, we obtain guaranteed coverage, however with larger $n_{cal}$ we obtain marginally better coverage. Irrespective of the chosen size of the calibration dataset, we obtain guaranteed coverage.
Figure 31: Impact of the size of the calibration dataset on the coverage obtained for Conformal Prediction across various nonconformity scores for the Wave Equation modelled using U-Net. Irrespective of the chosen size of the calibration dataset, we obtain guaranteed coverage.
Figure 32: Impact of the size of the calibration dataset on the coverage obtained for Conformal Prediction across various nonconformity scores for the Wave Equation modelled using FNO. Irrespective of the chosen size of the calibration dataset, we obtain guaranteed coverage.
From Figure 29 to Figure 32, we explore the impact of the size of the calibration dataset on the coverage obtained for the various nonconformity scores. We notice that across our experiments, we obtain guaranteed coverage, irrespective of the chosen nonconformity score or the size of the calibration dataset ($n_{cal}$). Though the size of the calibration dataset has no impact on the guarantee as can be witnessed in the above figures, the tightness of the error bars is governed by the size of the calibration dataset.